OpenStudy (anonymous):

solve for x 7(x-3)=4(x+5) a. x= 1/11 b. x= 1/3 c. x= 41/11 d. x= 41/3

OpenStudy (anonymous):

7(x-3) = 4(x+5) 7x - 21 = 4x + 20 3x = 41 x = 41/3 The answer is D. 41/3

OpenStudy (anonymous):

Thank you!